# Color code decoder with improved scaling for correcting circuit-level noise

### v1

@@ 4.1 Circuit and detectors
must be $+1$ when there are no errors. We refer to $L_{Z}$ as the logical observable of our scenario. After decoding, a correction $L_{Z}^{(\mathrm{corr})}$ is obtained and the decoding succeeds if and only if $L_{Z}=L_{Z}^{(\mathrm{corr})}$.
@@ 4.2 Generalization of the concatenated MWPM decoder
The Stim library [32] can be used to extract a DEM from a noise Clifford circuit (which is composed of Clifford gates, Pauli initializations/measurements, and Pauli noise channels) provided that detectors and logical observables are annotated appropriately. It works as follows: If the circuit has a single-Pauli error channel $\mathcal{E}_{q}^{(P)}$, which applies a specific Pauli product operator $P$ with probability $q$, we can get its effect by commuting $P$ to the end of the circuit and checking the detectors and logical observables flipped by it. If the circuit has a depolarizing channel $\mathcal{E}_{p}^{(1)}$ or $\mathcal{E}_{p}^{(2)}$, we can convert it into a sequence of single-Pauli error channels as
@@ 4             $\mathcal{D}_{P}\leftarrow\quantity{\text{$P$-type detectors in $\mathcal{D}$}}$;
5             $\mathcal{O}_{P}\leftarrow\mathcal{O}\text{ {if} }P=Z\text{ {else} }\emptyset$;
@@ 7                   Add $(q,\mathcal{D}_{P},\mathcal{O}_{P})$ to $\mathcal{M}_{ZX}$;
13Compress $\mathcal{M}_{ZX}$ (i.e., successively merge a pair of error mechanisms of $\mathcal{M}_{ZX}$ that have the same targets while updating the probability as $q=q_{1}+q_{2}-2q_{1}q_{2}$, where $q_{1}$ and $q_{2}$ are respectively the probabilities of the two error mechanisms);
@@ 18             Add $(q,\mathcal{D}_{\mathrm{\neg\texttt{c}}},\emptyset)$ to $\mathcal{M}_{\neg\texttt{c}}$;
22Compress $\mathcal{M}_{\neg\texttt{c}}$;
@@ 28             Add $(q,\mathcal{D},\mathcal{O})$ to $\mathcal{M}_{\texttt{c}}$;
Figure 6: Performance comparison of 292 length-7 CNOT schedules. (a) The failure rate $p_{\mathrm{fail}}$ of $T=7$ rounds of the logical idling gate when using the concatenated MWPM decoder is evaluated for code distance $d=7$ under the circuit-level noise model of strength $p=10^{-3}$ for each of these cnot schedules (sorted by the values of $p_{\mathrm{fail}}$ in ascending order). The error bars correspond to the 99% CIs. The selected optimal schedule is $[2,3,6,5,4,1;3,4,7,6,5,2]$, which is marked as a red dot. (b) The $Z$- and $X$-failure rates $p_{\mathrm{fail}}^{(Z)}$, $p_{\mathrm{fail}}^{(X)}$ for these cnot schedules are visualized as a scatter plot. The dashed line plots $p_{\mathrm{fail}}^{(X)}=p_{\mathrm{fail}}^{(Z)}$ and the gray area indicates the region of |$\absolutevalue{R_{\mathrm{bias}}}\leq 0.02$. The selected optimal schedule is marked as a red dot.
@@ 4.3 Optimization of the CNOT schedule
We note that our simulating scenario described in Sec. 2.3 is only for computing the $\overline{Z}$-failure rate $p_{\mathrm{fail}}^{(Z)}$ (i.e., failure rate of the $Z$ observable). Nonetheless, we can utilize the fact that the $\overline{X}$-failure rate $p_{\mathrm{fail}}^{(X)}$ is equal to the $\overline{Z}$-failure rate when using the cnot schedule with the “$Z$-part” (i.e., first six integers) and “$X$-part” (i.e., last six integers) reversed from the original one. For example, the $\overline{X}$-failure rate for the schedule $[2,3,6,5,4,1;3,4,7,6,5,2]$ can be computed from our scenario by instead using the schedule $[3,4,7,6,5,2;2,3,6,5,4,1]$. For a given cnot schedule, we compute the logical failure rate as
$p_{\mathrm{fail}}=p_{\mathrm{fail}}^{(X)}+p_{\mathrm{fail}}^{(Z)}-p_{\mathrm{fail}}^{(X)}p_{\mathrm{fail}}^{(Z)}$
by assuming that $\overline{X}$ and $\overline{Z}$ fail independently and quantify the bias between them as
@@
which is marked as red dots in Fig. 6. The corresponding failure rates and bias (99% CI) are
@@
Note that the failure rate for the worst-performing cnot schedule is $p_{\mathrm{fail}}=(2.9\pm 0.1)\times 10^{-3}$, which is approximately twice that of the best case.
Although we here choose the schedule to have a sufficiently small bias, biased schedules may be intentionally used depending on the task. For instance, ancillary logical qubits for the 15-to-1 magic state distillation scheme [35] are more detrimental to $\overline{X}$ errors than to $\overline{Z}$ errors as $\overline{Z}$ errors are detectable by the final $\overline{X}$ measurements [36], thus we can benefit from using cnot schedules biased to $\overline{Z}$ errors (i.e., $R_{\mathrm{bias}}>0$) for these qubits.
@@ 4.4 Performance analysis
Figure 7: Numerical analysis of the decoder under near-threshold circuit-level noise. We consider $T\in\quantity{1,2,4,8,16,32}$ rounds of the logical idling gate of the triangular color code with code distance $d\in\quantity{3,5,7,\cdots,29}$ under the circuit-level noise model of strength $p$. (a) Logical failure rates $p_{\mathrm{fail}}$ are plotted over $p$ for each $T$ and $d$. Each data point (marked as ‘X’) has the 99% CI of ±$\pm 10^{-3}$. For each $d$, the LOWESS regression of $p_{\mathrm{fail}}$ with fraction $2/3$ is drawn as a solid line. For each $d_{1}\in\quantity{13,17,21,25,29}$, the crossing of the two regression lines of $d=d_{1}$ and $d=\quantity(d_{1}+1)/2$ is highlighted in dark color. (b) The values of $p$ at these crossings $\widetilde{p}^{\times}_{\mathrm{circuit}}(d_{1},T)$ are plotted over $1/d_{1}$ for each $T$ with their linear regressions. The cross thresholds $p^{\times}_{\mathrm{circuit}}(T)$ are estimated by the intersections of these regressions with the vertical axis. (c) The 99% CIs of the estimations of $p^{\times}_{\mathrm{circuit}}(T)$ are presented with a fitted curve in the form of Eq. (2), where $p^{\times,\mathrm{LT}}_{\mathrm{circuit}}\approx 0.456\%$ and $\gamma\approx 1.02$. The orange dashed line plots $p^{\times}_{\mathrm{circuit}}(T)=p^{\times,\mathrm{LT}}_{\mathrm{circuit}}$.
@@
where $A$ is a free parameter. The values of $\widetilde{p}^{\times}_{\mathrm{circuit}}(d_{1},T)$ for $d_{1}\in\quantity{13,17,21,25,29}$ are highlighted in dark color in Fig. 7(a) and plotted in Fig. 7(b) against $1/d_{1}$ with their linear regressions. The obtained cross thresholds $p^{\times}_{\mathrm{circuit}}(T)$ are visualized in Fig. 7(c). By fitting them into Eq. (2), we get the long-term cross threshold $p^{\times,\mathrm{LT}}_{\mathrm{circuit}}\approx 0.456\%$ with $\gamma\approx 1.02$.
@@
$\begin{split}&p^{\star}_{\mathrm{circuit}}\approx 3.2\times 10^{-3},\quad\alpha\approx 9.3\times 10^{-3},\\ &\beta\approx 0.50,\quad\eta\approx 5.4,\quad d_{0}=12.\end{split}$
Figure 9: Comparison of three matching-based decoders under circuit-level noise. Logical failure rates per round $p_{\mathrm{fail}}/T$ are estimated by using three different decoders (projection, Möbius, and concatenated MWPM decoders) across different code distances $d$ at varying noise strengths $p\in\quantity{10^{-3},5\times 10^{-4},10^{-4}}$. The data for the projection decoder are from Fig. 17 of Ref. [14]. The data for the Möbius decoder are from Figs. 11 and 12 of Ref. [18], which respectively correspond to the superdense and middle-out color code circuits (two types of syndrome extraction circuits proposed in Ref. [18]). The solid lines represent the linear regressions of $\log p_{\mathrm{fail}}$ against $d$.
@@
In Figs. 1416 of Appendix C, we present separate analyses of the $\overline{Z}$- and $\overline{X}$-failure rates of the decoder, together with analysis of the bias $R_{\mathrm{bias}}$ defined in Eq. (8). To summarize, this is as follows: The long-term cross thresholds are estimated as $0.460\%$ for the $\overline{Z}$-failure and $0.452\%$ for the $\overline{X}$-failure, which differ by about 1.8%. The 99% CIs of the estimations of $R_{\mathrm{bias}}$ almost always include zero except for $d=3$, which implies that we cannot reject the null hypothesis of $R_{\mathrm{bias}}=0$ at the 99% significance level for these cases.
@@ Acknowledgements
We thank Felix Thomsen, Nicholas Fazio, Sam Smith, Benjamin J. Brown, and Michael E. Beverland for helpful discussions and comments. This work is supported by the Australian Research Council via the Centre of Excellence in Engineered Quantum Systems (EQUS) project number CE170100009. This article is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001122C0063. Any opinions, findings and conclusions or recommendations expressed in this article are those of the author(s) and do not necessarily reflect the views of the Defense Advanced Research Projects Agency (DARPA).
@@ A.2 Proof of validity when no boundaries
Using the projected boundary maps in Eq. (A.1), we rewrite Eq. (11b) as
@@ A.4 Superiority of the concatenated MWPM decoder over the projection decoder
Appendix B Discussion on small-weight uncorrectable errors
@@
B.1 Small-weight errors uncorrectable via the projection decoder but correctable via the concatenated MWPM decoder
@@
B.2 Small-weight errors uncorrectable via the concatenated MWPM decoder
Alike the Möbius MWPM decoder [17], the concatenated MWPM decoder fails to overcome the limitation that $O(3d/7)$-weight errors may not be correctable. We consider an error with weight $w=O(\xi d)$ for $\xi<1/2$ as presented in Fig. 11(a), which is composed of three string operators (one for each color) that terminate at the corresponding boundaries and make one check violated for each color. These string operators have the weights of $w_{\texttt{r}}=O(\xi_{\texttt{r}}d)$, $w_{\texttt{g}}=O([(\xi-\xi_{\texttt{r}})/2]d)$, and $w_{\texttt{b}}=w-w_{\texttt{r}}-w_{\texttt{g}}=O([(\xi-\xi_{\texttt{r}})/2]d)$, where $\xi_{\texttt{r}}\leq\xi$. Additionally, they are arranged in a way that each pair of violated checks (say, green and blue) are separated by $w_{\mathrm{gap}}=O([(1-\xi)/2]d)$ edges of the third color (red); that is, to move from the violated green check to the blue check, we need to pass at least $w_{\mathrm{gap}}$ red edges. Note that, if we move the violated blue and green checks to the violated red check and fuse them, it becomes a non-trivial string-net operator with weight $d$. When this error occurs, we face a problem that a first-round MWPM of a sub-decoding procedure can be wrong; for example, as shown in Fig. 11(b), the MWPM on the red-restricted lattice gives a wrong matching connecting the violated green and blue checks (drawn as a dotted purple line) if and only if

### v2

@@ 4.1 Circuit and detectors
must be $+1$ when there are no errors. We refer to $L_{Z}$ as the logical observable of our scenario. After decoding, a correction $L_{Z}^{(\mathrm{corr})}$ is obtained and the decoding succeeds if and only if $L_{Z}=L_{Z}^{(\mathrm{corr})}$. We describe $L_{Z}$ as ‘$Z$-type’ because it involves only $Z$-basis measurements and shares common factors ($\quantity{\widetilde{m}_{v}}$) exclusively with $Z$-type detectors; see Eqs. (6) and (7). Note that, for an alternative scenario where the logical qubit is initialized and measured in the $\overline{X}$ basis, we have one $X$-type logical observable instead.
@@ 4.2 Generalization of the concatenated MWPM decoder
The Stim library [38] can be used to extract a DEM from a noisy Clifford circuit (which is composed of Clifford gates, Pauli initializations/measurements, and Pauli noise channels) provided that detectors and logical observables are annotated appropriately. It works as follows: If the circuit has a single-Pauli error channel $\mathcal{E}_{q}^{(P)}$, which applies a specific Pauli product operator $P$ with probability $q$, we can get its effect by commuting $P$ to the end of the circuit and checking the detectors and logical observables flipped by it. If the circuit has a depolarizing channel $\mathcal{E}_{p}^{(1)}$ or $\mathcal{E}_{p}^{(2)}$, we can convert it into a sequence of single-Pauli error channels as
@@ 4             $\mathcal{D}_{P}\leftarrow\quantity{\text{$P$-type detectors in $\mathcal{D}$}}$;
             $\mathcal{O}_{P}\leftarrow\quantity{\text{$P$-type logical observables in $\mathcal{D}$}}$;
              // Our scenario has only one $Z$-type logical observable, but this is to ensure generalizability.
@@ 6                   Add $(q,\mathcal{D}_{P},\mathcal{O}_{P})$ to $\mathcal{M}_{ZX}$;
12Compress $\mathcal{M}_{ZX}$ (i.e., successively merge a pair of error mechanisms of $\mathcal{M}_{ZX}$ that have the same targets while updating the probability as $q=q_{1}+q_{2}-2q_{1}q_{2}$, where $q_{1}$ and $q_{2}$ are respectively the probabilities of the two error mechanisms);
@@ 17             Add $(q,\mathcal{D}_{\mathrm{\neg\texttt{c}}},\emptyset)$ to $\mathcal{M}_{\neg\texttt{c}}$;
21Compress $\mathcal{M}_{\neg\texttt{c}}$;
@@ 4.3 Optimization of the CNOT schedule
Figure 6: Performance comparison of 292 length-7 CNOT schedules. (a) The failure rate $p_{\mathrm{fail}}$ of $T=7$ rounds of the logical idling gate when using the concatenated MWPM decoder is evaluated for code distance $d=7$ under the circuit-level noise model of strength $p=10^{-3}$ for each of these cnot schedules (sorted by the values of $p_{\mathrm{fail}}$ in ascending order). The error bars correspond to the 99% CIs. The selected optimal schedule is $[2,3,6,5,4,1;3,4,7,6,5,2]$, which is marked as a red dot. (b) The $Z$- and $X$-failure rates $p_{\mathrm{fail}}^{(Z)}$, $p_{\mathrm{fail}}^{(X)}$ for these cnot schedules are visualized as a scatter plot. The dashed line plots $p_{\mathrm{fail}}^{(X)}=p_{\mathrm{fail}}^{(Z)}$ and the gray area indicates the region of |$\absolutevalue{R_{\mathrm{bias}}}\leq 0.02$. The selected optimal schedule is marked as a red dot. (c) Selected optimal schedule following the same graphical representation as in Fig. 5(a).
@@
We note that our simulating scenario described in Sec. 2.3 is only for computing the $\overline{Z}$-failure rate $p_{\mathrm{fail}}^{(Z)}$ (i.e., failure rate of the $Z$ observable). Nonetheless, we can utilize the fact that the $\overline{X}$-failure rate $p_{\mathrm{fail}}^{(X)}$ is equal to the $\overline{Z}$-failure rate when using the cnot schedule with the “$Z$-part” (i.e., first six integers) and “$X$-part” (i.e., last six integers) reversed from the original one. For example, the $\overline{X}$-failure rate for the schedule $[2,3,6,5,4,1;3,4,7,6,5,2]$ can be computed from our scenario by instead using the schedule $[3,4,7,6,5,2;2,3,6,5,4,1]$. For a given cnot schedule, we compute the logical failure rate as $p_{\mathrm{fail}}=p_{\mathrm{fail}}^{(X)}+p_{\mathrm{fail}}^{(Z)}$ and quantify the bias between them as
@@
Figure 7: Numerical analysis of the decoder under near-threshold circuit-level noise. We consider $T\in\quantity{1,2,4,8,16,32}$ rounds of the logical idling gate of the triangular color code with code distance $d\in\quantity{3,5,7,\cdots,29}$ under the circuit-level noise model of strength $p$. (a) Logical failure rates $p_{\mathrm{fail}}=p_{\mathrm{fail}}^{(X)}+p_{\mathrm{fail}}^{(Z)}$ are plotted over $p$ for each $T$ and $d$. Each data point (marked as ‘X’) has the 99% CI of ±$\pm 10^{-3}$. For each $d$, the LOWESS regression of $p_{\mathrm{fail}}$ with fraction $2/3$ is drawn as a solid line. For each $d_{1}\in\quantity{13,17,21,25,29}$, the crossing of the two regression lines of $d=d_{1}$ and $d=\quantity(d_{1}+1)/2$ is highlighted in dark color. (b) The values of $p$ at these crossings $\widetilde{p}^{\times}_{\mathrm{circuit}}(d_{1},T)$ are plotted over $1/d_{1}$ for each $T$ with their linear regressions. The cross thresholds $p^{\times}_{\mathrm{circuit}}(T)$ are estimated by the intersections of these regressions with the vertical axis. (c) The 99% CIs of the estimations of $p^{\times}_{\mathrm{circuit}}(T)$ are presented with a fitted curve in the form of Eq. (2), where $p^{\times,\mathrm{LT}}_{\mathrm{circuit}}\approx 0.455\%$ and $\gamma\approx 1.01$. The orange dashed line plots $p^{\times}_{\mathrm{circuit}}(T)=p^{\times,\mathrm{LT}}_{\mathrm{circuit}}$.
@@
which is marked as red dots in Figs. 6(a) and (b) and visualized in Fig. 6(c). The corresponding failure rates and bias (99% CI) are
@@
Note that the failure rate for the worst-performing cnot schedule is $p_{\mathrm{fail}}=(2.9\pm 0.1)\times 10^{-3}$, which is approximately twice that of the best case. In Appendix B, we discuss the origin of the bias and explain why the degree of this bias varies with the cnot schedule.
Although we here choose the schedule to have a sufficiently small bias, biased schedules may be intentionally used depending on the task. For instance, in the 15-to-1 magic state distillation scheme [49], $\overline{X}$ errors on ancillary logical qubits are more detrimental than $\overline{Z}$ errors, as the latter can be detected by the final $\overline{X}$ measurements [43]. Therefore, using cnot schedules biased towards $\overline{Z}$ errors (i.e., $R_{\mathrm{bias}}>0$) for the ancillary logical qubits can be advantageous.
@@ 4.4 Performance analysis
where $A$ is a free parameter. The values of $\widetilde{p}^{\times}_{\mathrm{circuit}}(d_{1},T)$ for $d_{1}\in\quantity{13,17,21,25,29}$ are highlighted in dark color in Fig. 7(a) and plotted in Fig. 7(b) against $1/d_{1}$ with their linear regressions. The obtained cross thresholds $p^{\times}_{\mathrm{circuit}}(T)$ are visualized in Fig. 7(c). By fitting them into Eq. (2), we get the long-term cross threshold $p^{\times,\mathrm{LT}}_{\mathrm{circuit}}\approx 0.455\%$ with $\gamma\approx 1.01$.
@@
$\begin{split}&p^{\star}_{\mathrm{circuit}}\approx 3.2\times 10^{-3},\quad\alpha\approx 8.4\times 10^{-3},\\ &\beta\approx 0.50,\quad\eta\approx 5.4,\quad d_{0}=12.\end{split}$
Figure 9: Comparison of three matching-based decoders under circuit-level noise. Logical failure rates per round $p_{\mathrm{fail}}/T$ are estimated by using three different decoders (projection, Möbius, and concatenated MWPM decoders) across different code distances $d$ at varying noise strengths $p\in\quantity{10^{-3},5\times 10^{-4},10^{-4}}$. The data for the projection decoder are from Fig. 17 of Ref. [20]. The data for the Möbius decoder are from Figs. 11 and 12 of Ref. [24], which respectively correspond to the superdense and middle-out color code circuits (two types of syndrome extraction circuits proposed in Ref. [24]). The solid lines represent the linear regressions of $\log(p_{\mathrm{fail}}/T)$ against $d$.
@@
In Figs. 1517 of Appendix D, we present separate analyses of the $\overline{Z}$- and $\overline{X}$-failure rates of the decoder, together with analysis of the bias $R_{\mathrm{bias}}$ defined in Eq. (8), while setting the cnot schedule to Eq. (9) as well. The results are summarized as follows: The long-term cross thresholds are estimated as $0.460\%$ for the $\overline{Z}$-failure and $0.452\%$ for the $\overline{X}$-failure, which differ by about 1.8% (0.008%p). However, this difference likely falls within the range of statistical error, as the 99% CIs for the estimations of $R_{\mathrm{bias}}$ almost always include zero, implying that the null hypothesis of $R_{\mathrm{bias}}=0$ cannot be rejected at the 99% significance level for these cases.
@@ 5 Remarks
Another promising direction for future research is to apply our concatenation approach to the Möbius decoder [23, 24]. Specifically, instead of considering the three colors independently, we could run the first-round MWPM on the unified lattice, as done in the Möbius decoder. A key challenge lies in constructing the lattice for the second-round MWPM and lifting the first-round MWPM outcomes from the Möbius decoder to this lattice. We have two preliminary ideas for this: We could either (i) combine the three monochromatic lattices by introducing extra edges to account for errors at the boundaries between different colors, or (ii) keep each monochromatic lattice independent and lift each boundary error to two nodes belonging to distinct monochromatic lattices. Developing and testing these ideas further would be a worthwhile direction for future work.
@@ Acknowledgements
We thank Felix Thomsen, Nicholas Fazio, Samuel C. Smith, Benjamin J. Brown, Michael E. Beverland, and Kaavya Sahay for helpful discussions and comments. This work is supported by the Australian Research Council via the Centre of Excellence in Engineered Quantum Systems (EQUS) project number CE170100009. This article is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001122C0063. Any opinions, findings and conclusions or recommendations expressed in this article are those of the author(s) and do not necessarily reflect the views of the Defense Advanced Research Projects Agency (DARPA).
@@ A.2 Proof of validity when no boundaries
Using the projected boundary maps in Eq. (10), we rewrite Eq. (11b) as
@@ A.4 Superiority of the concatenated MWPM decoder over the projection decoder
Appendix B Discussion on the origin of the bias between $\overline{X}$ and $\overline{Z}$ failures
In this appendix, we discuss the origin of the bias between $\overline{X}$ and $\overline{Z}$ failures and explain why the degree of this bias varies with the cnot schedule. Errors occurring during syndrome extraction propagate to other data qubits via cnot gates, and the propagation pattern strongly depends on the cnot schedule. As a result, it is not surprising that failure rates are influenced by the schedule. In particular, hook errors, which are errors on ancillary qubits that occur midway through syndrome extraction and propagate to more than one data qubit, can significantly impact the decoder’s performance, as they may reduce the fault distance.
To understand the origin of the bias, it is helpful to compare the optimal schedule,
$\mathcal{A}_{\mathrm{optimal}}=[2,3,6,5,4,1;3,4,7,6,5,2],$ (14)
which exhibits a very small bias (discussed in Sec. 4.3), with the schedule that produces the highest bias,
$\mathcal{A}_{\mathrm{biased}}=[1,6,7,5,4,2;2,3,6,7,5,4],$
where $R_{\mathrm{bias}}\approx 0.99$ and $p_{\mathrm{fail}}^{(X)}\approx 3p_{\mathrm{fail}}^{(Z)}$ for $d=T=7$, $p=10^{-3}$. The optimal schedule has the property that its $Z$-part and $X$-part follow the same pattern, differing consistently by 1. This symmetry ensures that hook errors affect data qubits in the same way for the two ancillary qubits on a single face. In contrast, the schedule $\mathcal{A}_{\mathrm{biased}}$ lacks such consistency; its $Z$- and $X$-parts are entirely different, resulting in hook errors exhibiting distinct characteristics for the two Pauli types.
Figure 10: Possible hook errors on the distance-3 color code patch with the CNOT schedule in Eq. (14), producing the highest bias. Each purple (orange) box labeled ‘$x$$y$’ adjacent to an edge between two qubits represents a $Z\otimes Z$ ($X\otimes X$) error on these qubits, which is caused by a Pauli-$Z$ ($X$) hook error occurring between time steps $x$ and $y$ on the nearby $Z$-type ($X$-type) ancillary qubit.
To illustrate this effect explicitly, Fig. 10 shows a distance-3 patch for schedule $\mathcal{A}_{\mathrm{biased}}$, with possible hook errors visualized using purple and orange boxes. For instance, the purple box adjacent to an edge connecting qubits A and B (labeled ‘2–4’ with ‘Z Z’ inside) represents a $Z_{\mathrm{A}}\otimes Z_{\mathrm{B}}$ error caused by a hook error between time steps 2 and 4 on the nearby $Z$-type ancillary qubit (located on face A-B-C-D). It is evident that $Z$-type hook errors are more likely to occur than $X$-type hook errors, as they involve a greater number of time steps (e.g., $Z_{\mathrm{F}}\otimes Z_{\mathrm{G}}$ involves time steps 2 to 6). Consequently, $\overline{Z}$ errors are expected to be more frequent than $\overline{X}$ errors, which aligns with the numerical results. Generalizing this argument to higher code distances might be slightly more complex due to weight-6 checks, but the basic principles remain the same.
Appendix C Discussion on small-weight uncorrectable errors
@@
C.1 Small-weight errors uncorrectable via the projection decoder but correctable via the concatenated MWPM decoder
@@
C.2 Small-weight errors uncorrectable via the concatenated MWPM decoder
Alike the Möbius MWPM decoder [23], the concatenated MWPM decoder fails to overcome the limitation that $O(3d/7)$-weight errors may not be correctable. We consider an error with weight $w=O(\xi d)$ for $\xi<1/2$ as presented in Fig. 12(a), which is composed of three string operators (one for each color) that terminate at the corresponding boundaries and make one check violated for each color. These string operators have the weights of $w_{\texttt{r}}=O(\xi_{\texttt{r}}d)$, $w_{\texttt{g}}=O([(\xi-\xi_{\texttt{r}})/2]d)$, and $w_{\texttt{b}}=w-w_{\texttt{r}}-w_{\texttt{g}}=O([(\xi-\xi_{\texttt{r}})/2]d)$, where $\xi_{\texttt{r}}\leq\xi$. Additionally, they are arranged in a way that each pair of violated checks (say, green and blue) are separated by $w_{\mathrm{gap}}=O([(1-\xi)/2]d)$ edges of the third color (red); that is, to move from the violated green check to the blue check, we need to pass at least $w_{\mathrm{gap}}$ red edges. (Note that a non-trivial string-net operator constructed by moving the violated blue and green checks to the violated red check and fusing them has a weight of $w+2w_{\mathrm{gap}}=O(d)$.) When this error occurs, we face a problem that a first-round MWPM of a sub-decoding procedure can be wrong. For example, as shown in Fig. 12(b), the MWPM on the red-restricted lattice, which involves only the violated green and blue checks, gives a wrong matching connecting these checks (drawn as a dotted purple line) if and only if the distance between them ($w_{\mathrm{gap}}$) is smaller than the sum of their distances to the corresponding boundaries ($w_{\texttt{g}}+w_{\texttt{b}}$) within the red-restricted lattice, i.e.,